Beginning three months from now, you want to be able to withdraw $2,300 each quarter from your bank account to cover college expenses over the next four years. If the account pays 1.45 percent interest per quarter, how much do you need to have in your bank account today to meet your expense needs over the next four year

A. 28,785.91

B. 32,633.32

C. 41,086.02

D. 19,934.56
Answer #1

I/Y = Rate = 1.45%

Nper = Number of periods = 4*4 = 16

PMT = Annual payment = -2300

FV = future value = 0

PV = ?

______________________________________

Excel function PV

=PV(1.45%,16,-2300,0)

Answer = 32,633.32

Note: Years is multiplied with 4. Reason is quarterly compounding.